A father is now 7 times as old as his son. In 5 years he will be 4 times as old as his son will be then. What are their present ages?
Answer by Boreal(15235)

son is x
father is 7x
7x+5=4(x+5) in 5 years
7x+5=4x+20
3x=15
x=5
son is 5, father is 35 ANSWER
in 5 years, son is 10, father is 40.